OpenStudy (anonymous):

Correct me if I'm wrong? a over 3 all over 6 over 3a ? I got 6a/6a

OpenStudy (ash2326):

\[\Large \frac{\frac a 3}{\frac 6 {3a}}\] we'll get \[\frac{a\times 3a}{3\times 6}\] can you do it now?

OpenStudy (anonymous):